4. The (n – 1)th term of an A.P. is given by 7,12,17, 22,… is
(a) 5n + 2
(b) 5n + 3
(c) 5n – 5
(d) 5n – 3

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: d Explaination:Reason: Here a = 7, d = 12-7 = 5 ∴ an-1 = a + [(n – 1) – l]d = 7 + [(n – 1) -1] (5) = 7 + (n – 2)5 = 7 + 5n – 10 = 5M – 3